OpenStudy (anonymous):

How to find missing length using pythagorean theorem and special right triangles

OpenStudy (e.mccormick):

If you think about it, the distance fomula is the pythagorean theorem. You just use the difference in x and difference in y as the things you square.

OpenStudy (e.mccormick):

As for "specal" right triangles, they are just some known ratios you an use easily.

OpenStudy (anonymous):

i dont know what to do with the square root numbers like 2sqrt2, 3sqrt2, and 5sqrt3 and it gives me two other variables

OpenStudy (e.mccormick):

Well, let me show you the simplest example: \(x^2=4\) So we take the square root of both sides. \(\sqrt{x^2}=\sqrt{4}\) and for a distance, you only use the primary positive root, so: \(x=2\) Now, with the Pythagorean Formula, you might get something more like: \(a=3\), \(b=4\), \(c=?\) when \(a^2+b^2=c^2\) So you plug it in: \(3^2+4^2=c^2\) \(9+16=c^2\) \(25=c^2\) \(\sqrt{25}={c^2}\) \(5=c\) And eventially they move to things that don't have an integer root, so the answer might be \(3\sqrt{2}\), which is still fine.

OpenStudy (triciaal):

pythagorean theorem a^2 + b^2 = c^2 where c is the hypotenuse, the longest side, and a and b are the legs just substitute and do the evaluation

OpenStudy (e.mccormick):

Numbers under roots are just real numbers that you can use. They don't necessarily have a nice, clean stop to them when you take them out from under the root.

OpenStudy (anonymous):

i have one square root(ones i listed) and two variables

OpenStudy (e.mccormick):

Without knowing where they are in the diagram or on the triangle, it is hard to say what to do with them.

OpenStudy (triciaal):

2 most common special right triangles are 3,4, 5 as used above and 5, 12, 13

OpenStudy (anonymous):

the other thing is angles like 45 degrees 30 degrees and 60 degrees

OpenStudy (e.mccormick):

All 30-60-90 have the same rations to the sides: http://www.themathpage.com/atrig/30-60-90-triangle.htm There is a similar set of inforation for 45-45-90

OpenStudy (triciaal):